OpenStudy (anonymous):

what is the equation of the line with the slope 5/2 and passes through (-2,1)

OpenStudy (tkhunny):

What does the equation of such a line look like? Perhaps you are familiar with the slope-intercept form?

OpenStudy (anonymous):

y = mx + b

OpenStudy (tkhunny):

Perfect. You are given the slope. Substitute that known value and rewrite the present state of the equation of the line.

OpenStudy (anonymous):

Heres what I did: y= 5/2 x + b 1 = 5/2 (-2/1) + b 1 = 10/2 + b 1 = 5 + b -4 = b y= 5/2x - 4

OpenStudy (tkhunny):

Okay, after the slope we have y = (5/2)x + b Substituting the known point gives: 1 = (5/2)(-2) + b ==> 1 = -5 + b ==> b = 6 Looks like you lost a sign in there, somewhere.

OpenStudy (anonymous):

ohhhhh, I caught on to my mistake ! I didn't realize the -2/1 when I wrote it! Sweet. I was on the right path. Awesome. Thanks.

OpenStudy (tkhunny):

It would help if you would seek ways to make your life simpler. You did this \(\dfrac{5}{2}\cdot(-2) = \dfrac{5}{2}\cdot\dfrac{-2}{1} = \dfrac{-10}{2} = -5\) That's a lot of trouble. No Wonder you lost a sign on the way. I did this: \(\dfrac{5}{2}\cdot(-2) = 5\cdot\dfrac{-2}{2} = 5(-1) = -5\) So much less to go wrong! You have to train yourself to look for such opportunities. Don't get stuck in one way of thinking or solving. There may be many ways to go about it.

OpenStudy (anonymous):

Wow. That was super simple. Lol. I will be training myself from this point on to use that method. That was way quicker than how I actually did it.